RC meeting and Run Plan, Thursday January 19th
Return to the RC Meeting Calendar
RC Meeting Information
|Current RC:||Iuliia Skorodumina|
|Time:||Daily at 3:30PM|
|Location:||Counting House Conference Room|
Remote Connection Details
- Join ZoomGov Meeting
- Meeting ID: 161 903 2184
- Passcode: 260715
- Dial by your location
- +1 669 254 5252 US (San Jose)
- +1 646 828 7666 US (New York)
- +1 551 285 1373 US
- +1 669 216 1590 US (San Jose)
- 833 568 8864 US Toll-free
- Find your local number: https://jlab-org.zoomgov.com/u/acj3pTicMx
We started Wednesday Swing Shift (01/18) in Controlled Access.
GEM experts were working in the hall (see https://logbooks.jlab.org/entry/4112604). GEMs need further software alignment. GEM experts will do the alignment once we start taking data.
Jiwan was working on HCal. See his logentry https://logbooks.jlab.org/entry/4112681
Around 19:30 the Hall went to the Beam Permit. MCC started beam tuning around 8pm.
Around 22:30 the CW beam was sent and we started the Beam Checkout. Bill provided detailed instructions to the Shift Crew (see https://logbooks.jlab.org/entry/4112725) and was monitoring the Beam Checkout remotely.
The beam turned out to be very wide in X. MCC was trying to reduce the beam size up until 5am, but the beam remained too wide. At 5 am the beam was gone and did not resume until the end of the Owl shift (8am). See Owl shift summary (https://logbooks.jlab.org/entry/4112938).
Since we did not succeed with establishing production beam during Owl, we cannot take 45 uA on glass cells during the MCC current push. For this reason, MCC suggested us to take up to 45 uA on a Carbon target (Carbon Hole) for the purpose of the current push and to establish the production beam afterwards. However, we cannot proceed with this plan until we check the raster size and update operational restrictions accordingly (as recently we made changes to the raster coils/generators). The MCC current push was postponed until 10 am.
Thursday Day Shift (01/19). From 8 am to 9:30 am Jiwan was working on checking the raster size on a Carbon Hole. Once the raster was checked, we contacted Silviu and asked him to update operational restrictions. After this we were ready to take up to 45 uA on Carbon Hole. The MCC started the current push around 10 am.
The high current push was in action the whole Day Shift and was successful. Now 45 uA can be delivered to Hall A. The Summary of the high current push: https://logbooks.jlab.org/entry/4113138 . MCC is planning on finishing the current push at around 5pm and once they are done we plan to resume establishing production beam to the hall.
Tomorrow (Friday, 01/20) MCC plans the accelerator repair work (see PD Update https://logbooks.jlab.org/entry/4113148). The work is scheduled to start at 7:15am and continue up to the end of the Day Shift. The hall will go to the Controlled Access.
Work in the hall during the Controlled Access (01/20)
- GEM work, which include plugging in the loose cable https://logbooks.jlab.org/entry/4113050
- HCAL work. Jiwan will continue troubleshooting noisy channels https://logbooks.jlab.org/entry/4112681
- Timing Hodoscope problematic channel (Ralph Marinaro)
- anything else?
Please consider signing up for shifts .